Im not sure what topic it would fall under. But here is my questions
the variable x is 4 less than the variable y?
answer I think it is, x - 4 = y
next question
if the variable t is multiplied by 5, the result is equal to the variable r.
answer I think it is, t * 5 = r
next question
if 3 is subtracted from the product of the variables x and y the result is equal to 4 time the variable z
this one I really couldn't understand
next question
the quotient of the variable x with the variable y is 2 more than o.7 of the variable z.
this is another one I don't understand.
If there is a way to explain the answer to me and how you get the answer I would appreciate it.

Answer by venugopalramana(3286)

Im not sure what topic it would fall under. But here is my questions
the variable x is 4 less than the variable y?
THINK THIS WAY..WHICH IS LESS OR MORE..X IS LESS..Y IS MORE..SO TO GET SMALLER NUMBER X FROM A BIGGER NUMBER Y YOU NEED TO SUBTRACT 4 FROM BIGGER NUMBER.HENCE ANSWER IS X=Y-4
answer I think it is, x - 4 = y...NO...Y-4=X
next question
if the variable t is multiplied by 5, the result is equal to the variable r.
answer I think it is, t * 5 = r...OK..GOOD
next question
if 3 is subtracted from the product of the variables x and y the result is equal to 4 time the variable z
GO STEP BY STEP
PRODUCT OF X AND Y =XY
3 IS SUBTRACTE..FROM THE PRODUCT ..SO WE GET
XY-3
4 TIMES Z =4Z
BOTH ARE SAME..SO
XY-3=4Z
this one I really couldn't understand
next question
the quotient of the variable x with the variable y is 2 more than o.7 of the variable z.
this is another one I don't understand.